core: 0.67 of 1 pt 17 of 20 (20 complete) ▼ HW Score: 98.33%, 1967 of 20 pts Problem 1.14 Question Help * Charles Lackey operates a bakery in Idaho Falls, Idaho. Because of its excellent product and excellent location, demand has increased by 45% in the last year. On far too many occasions, customers have not been able to purchase the bread of their choice. Because of the size of the store, no new ovens can be added. At a staff meeting, one employee suggested ways to load the ovens differently so that more loaves of bread can be baked at one time. This new process will require that the ovens be loaded by hand, requiring additional manpower. This is the only production change that will be made in order to meet the increased demand. The bakery currently makes 1,500 loaves per month. Employees are paid S8 per hour. In addition to the labor cost, Charles also has a constant utility cost per month of $800 and a per loaf ingredient cost of SO.40 Current multifactor productivity for 640 work hours per month = 0.230 loaves/dollar (round your response to three decimal places) After increasing the number of work hours to 928 per month, the multifactor productivity loaves/dollar (round your response to three decimal places)Please help I dont understand how to get the second part!
Answer #1

Multi factor productivity= Total output / ( total labor+ total COST + total ingredient cost)

Total output=1500

Total labor=$8*928=$7424

Utility cost=$800

Ingredient cost= $0.4*1500=$600

Total cost = $7424+$800+$600= 8824

Multi factor productivity= Total output / ( total labor+ total COST + total ingredient cost)

Multi factor productivity= 1500/8824=0.16999loaves / dollar

KINDLY RATE THE ANSWER AS THUMBS UP. THANKS, A LOT.